OpenStudy (anonymous):

what is the solution to the system of equations? x+y=5 x-2y=2

OpenStudy (anonymous):

i think it is 4,1

OpenStudy (anonymous):

from the first equation we have: x = 5 - y, now plug this expression of x in the second equation: (5-y) - 2y = 2 5 - 3y = 2 y = 1 Now just find x

OpenStudy (anonymous):

find x here: x = 5 - y

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

No problem ;)

OpenStudy (anonymous):

can u help me with 1 more

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

x=3y+5 x-4y=3

OpenStudy (anonymous):

from the first equation, we already have: x = 3y + 5, plug thsi expression of x inside the second equation: x - 4y = 3 (3y + 5) - 4y = 3 -y + 5 = 3 y = 2, now find x

OpenStudy (anonymous):

11?

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

yeah!!! thanks so much

OpenStudy (anonymous):

No problem ;)